OpenStudy (anonymous):

Solve x2 − 8x − 20 = 0. x = 4, x = −5 x = −4, x = 5 x = 2, x = −10 x = −2, x = 10

Parth (parthkohli):

I suggest that solving by factoring is a good idea here.

Parth (parthkohli):

A good start to solving this particular problem is of course — factoring. We start it by: \(\Rightarrow x^2 - 10x + 2x + 20 = 0 \)

Parth (parthkohli):

Then group it. \(\Rightarrow x(x - 10) + 2(x - 10) = 0\) \(\Rightarrow (x + 2)( x - 10) = 0\)